OpenStudy (anonymous):

the speed of a passenger train is 6 mph faster than the speed of a freight train. the passenger train travels 300 miles in the same time it take the freight train to travel 270 miles. find the speed of each train

OpenStudy (saifoo.khan):

I hate this one! >_<

OpenStudy (saifoo.khan):

Let me solve please.

OpenStudy (anonymous):

\[t=\frac{d}{r} \]Let x = the speed of the freight train. Solve the following for x.\[\frac{300}{x+6}= \frac{270}{x}\]x = 54mph, x+6 = 60 mph